# Desynchronization Index: A New Connectivity Approach for Exploring Epileptogenic Networks

### v3

@@
Abstract
Objective: In this work, we propose a new computational framework to assist neurophysiologists in Stereoelectroencephalography (SEEG) analysis, with the final aim of improving the definition of the Epileptogenic Zone (EZ) in patients with drug-resistant epilepsy. Method: We consider the Phase Transfer Entropy (PTE) model to estimate the effective connectivity between SEEG channels. Hence, we design an algorithm, named the Desynchronization Index (DI), that classifies as epileptogenic the channels that show independent behavior with respect to the rest of the network during the seconds preceding the seizure propagation. Results: We test the proposed DI algorithm against Epileptogenic Index (EI) on a clinical dataset of 11 patients, considering the neurophysiological evaluation of the EZ as the clinical ground truth. Our results denote that DI overcomes EI in terms of area under the ROC curve (AUC=0.80 vs AUC=0.74), while combining the two algorithms as a unique tool leads to the best performance (AUC=0.87). Conclusions: The DI algorithm underscores connectivity dynamics that can hardly be identified with a pure visual analysis, increasing the accuracy in the EZ definition compared to traditional methods. Significance: The integration of connectivity- and energy-based metrics, as the one we propose, can lead to the definition of a new effective biomarker of the EZ, reducing the burden required by the SEEG review in the case of extensive implants and improving our understanding of the dynamics leading to the generation of seizures.
{IEEEkeywords}
@@ Abstract
This paper has been submitted to IEEE Transactions on Biomedical Engineering
1 Introduction
[…]
@@ 1 Introduction
The localization of the Epileptogenic Zone (EZ) is the fundamental prerequisite for performing epilepsy surgery in people with drug-resistant epilepsy [1]. Ensuring a higher accuracy in the definition of the EZ makes it possible to reduce the extent of the surgery, mitigating the risk of complications or chronic deficits. In the most complex scenario, the EZ can be identified only through invasive procedures, among which Stereoelectroencephalography (SEEG) represents the most accurate methodology [2]. An SEEG recording involves the surgical implantation of electrodes into the patient’s brain, enabling the monitoring of electrical activity from deep cortical regions, which is essential for the identification of the EZ. As a result, SEEG-guided surgical resections lead more than $60\%$ of patients to be seizure-free after the intervention [3]. However, SEEG has the drawback of providing very focused information: if none of the intracranial electrodes intersects the EZ, the procedure may result in an unclear EZ definition and an unsuccessful epilepsy surgery [4].
@@
Nowadays, SEEG interpretation is mainly based on visual analysis, focusing on the early phase of the seizure discharge. The final goal is to define an EZ that can be focal or distributed among multiple cortical structures, a scenario falling under the concept of epileptogenic networks [5]. The above process is particularly complex since SEEG signals include hundreds of components, whose analysis requires the involvement of highly specialized neurophysiologists. For this reason, in the last years, SEEG interpretation has started being supported by computational tools apt at characterizing the EZ and, more in general, the dynamics leading to seizure generation [6]. The most common methods analyze the SEEG power spectrum, searching for the cortical sites generating fast oscillations [7], i.e., electrical activities with frequencies in the gamma range ($[30,100]$ Hz), as done in the cases of Epileptogenic Index (EI) [8] and Epileptogenic Maps (EMs) [9].
@@
Other approaches try to model the cortical connectivity, identifying which brain structures show abnormal connectivity behavior during the seizure onset. In this context, the first studies have focused on the electric responses generated through intracranial stimulations, known as Cortico-Cortical Evoked Potentials (CCEPs) [10]. The SEEG connectivity can also be estimated by analyzing the phase distribution of the electrical signals generated by each cortical site, as performed by the Phase Locking Value (PLV) and the Phase Lag Index (PLI), avoiding biases associated with amplitude differences.
@@
Although multiple quantitative approaches have been developed for the study of epileptogenic networks, the literature often presents discordant findings. Most works agree that the ictal phase (i.e., the period involving the seizure discharge) is preceded by a connectivity reduction, followed by the opposite phenomenon during the seizure propagation [11]. Recently, it has been hypothesized that the EZ presents an abnormal inward information flow during the inter-ictal phase [12], and the epileptogenic level of each cortical site can be assessed by the connection exerted during the epileptic discharge [13].
@@
In this work, we present a new framework for supporting neurophysiologists in SEEG interpretation. Specifically, we design a connectivity-based algorithm, named Desynchronization Index (DI), that exploits the Phase Transfer Entropy (PTE) to estimate the relations between brain structures and detect connectivity anomalies in the resulting time-varying network. According to our approach, the EZ is identified by the cortical sites that present an independent behavior during the ictal phase. Overall, the main contributions of this manuscript consist of the following points:
@@ 2 Related Work
In clinical practice, the interpretation of SEEG recordings is mainly based on visual analysis, which is a highly time-consuming process because of the high number of signals that must be reviewed [14]. The most common marker of the seizure onset is given by low-voltage fast activities, i.e., electrical discharges spanning beta frequencies ($[13,30]$ Hz), usually observed in mesial temporal seizures, and gamma frequencies ($[30,100]$ Hz), observed in neocortical seizures. The quantification of fast activities and the network dynamics leading to seizures can be achieved by analyzing the SEEG spectrum. Such an approach is followed by the EI algorithm [8], which represents the most routinely used method to assist clinicians in SEEG interpretation [15].
@@
The EI algorithm, occasionally combined with other biomarkers [16], has been recognized to provide the best accuracy in terms of EZ localization and surgical prediction. In this regards, the epileptogenicity rank extends the EI framework by assigning weights to SEEG sites depending on to their distance from the hypothesized center of the EZ [17]. This technique has the obvious limitation of biasing the outcome of the quantitative analysis by a prior assumption. Hence, the epileptogenicity rank reduces the number of false positives when the clinical hypothesis is correct, but dramatically decrease the algorithm sensitivity in the other case.
@@
While EI focus on the energy distribution in the frequency domain, other approaches try to estimate the variation in neuronal connectivity during the ictal phase. Several tools have been proposed for such a goal, including the PLV [18] and the PLI [19], which involve different trade-offs in terms of sensitivity and precision [20]. In particular, the PLI has the advantage of avoiding the noisy information associated with the volume conduction [21], reducing the number of false positives. A more advanced method is the PTE [22], which exploits the Granger Casuality (GC) [23] to estimate both the delay and the magnitude of neuronal connections, representing a very promising technique for SEEG analysis [24].
@@
In addition to designing EI, Bartolomei et al. analyze the SEEG data using the nonlinear regression coefficient (h2) [25], and show that the epileptogenic network is associated with reduced connectivity in the period that anticipates the ictal phase  [26]. While the h2 coefficient is agnostic to the signal frequencies, most connectivity models discern cortical relationships depending on the frequency bands. Following this principle, the authors of [27] estimate connectivity via linear correlation and observe that seizure frequency is proportional to the EZ outward connectivity. Similar methodologies are given in [28], whose pairwise correlation is combined with Euclidean distance to estimate the EZ extension, and [29], where graph theory is exploited to identify the network hubs that most contribute to the seizure propagation.
@@
Recently, the use of Spearman rank correlation [30] to assess the coupling between brain structures has shown good results for both the detection of the EZ and the prediction of surgical outcomes [31]. Other connectivity-based approaches associate the phase of slow oscillations ($<30$ Hz) with the amplitude of fast oscillations ($>30$ Hz), as occurred for the Phase Slope Index (PSI) and the Phase Amplitude Coupling (PAC) [32]. Exploiting the PSI, [33] shows that the EZ presents higher outward connectivity during the seizure onset than the resting state. Similar results are observed in [34], where the PAC between high (gamma) and low (delta and theta) frequency bands is computed. The authors of [35] seem to contradict these findings and, using the Partial Direct Coherence (PDC) [36] as a connectivity measure, assess that the EZ presents higher inward connectivity during seizure generation.
@@
In recent years, the popularity of Machine Learning (ML) has encouraged many researchers to apply data-driven techniques to epileptogenic networks. For example, the authors of [37] exploit hypergraph learning to automatically detect SEEG fast oscillations, while [38] investigates different patient-specific ML models for analyzing scalp electroencephalography signals. In addition, [39] exploit a Support Vector Machine (SVM) model to integrate multiple electrophysiological characteristics, including both fast oscillations and low-frequency suppression, to improve accuracy in EZ detection. This supports the fact that ictal channels are not identified by specific frequency components, and more holistic methodologies are required to describe the generation of seizures. On the other hand, [39] and similar works consider as ground truth the area resected by epilepsy surgery, which is usually far larger than the EZ to maximize the probability of seizure reduction. In general, the high complexity behind SEEG analysis makes it extremely difficult to build supervised datasets, which prevents ML solutions from obtaining reliable performance in this field [40].
@@
Although the large variety of works implementing connectivity-based metrics for the interpretation of SEEG, the results of the literature do not lead to an agreement on which biomarkers characterize EZ. Initial works show that the ictal transition is characterized by a reduction in connectivity, but the EZ is often identified by cortical sites with a greater coupling during seizure propagation. In this work, we try to answer these questions by quantifying the level of epileptogenicity of SEEG sites in terms of desynchronization. With this latter, we denote the tendency of a cortical site to show independent electrical behavior that cannot be inhibited by neighboring brain structures.
3 Analytical Model
@@
In this section, we first describe how SEEG signals are modeled within our framework and introduce the EI algorithm, which represents the state-of-the-art for detecting the EZ. Then, we present the connectivity model used to infer the relation between SEEG sites starting from their phase distribution. Finally, we introduce the DI algorithm, which evaluates the epileptogenicity level of each site according to its tendency to desynchronize from the network.
3.1 Epileptogenic Index
@@
We model a single SEEG recording as a multidimensional signal with $N$ different components, named channels, one for each cortical site analyzed. In the following, we denote by $\mathcal{N}$ the set of channels and by ||$\cdot$|| ⋅ | the cardinality operator, so that the number of cortical sites is equal to $N=|\mathcal{N}|$. According to our approach, each channel $x\in\mathcal{N}$ is segmented in multiple overlapping windows $x(t)$, with $t=0,\Delta t,2\Delta t,…$, where $\Delta t$ represents the time-shift between consecutive windows. Given the window duration $T_{\text{window}}$ and the sampling frequency $f_{s}$, the sample number per window is $n=T_{\text{window}}/f_{s}$.
@@
The EI algorithm requires to evaluate the signal energy in the frequency domain [8]. Given a specific window $x(t)$, the algorithm first computes the Fourier Transform (FT) of $x(t)$, obtaining a complex value $X(t,f)$ for each frequency in $[0,f_{s}/2]$. Hence, the energy ratio between the high- and low-frequency bands of the target signal $x$ is computed as
@@
$\Gamma(\omega,t)=\begin{cases}\max\left\{0,\Gamma(\omega,t-\Delta t)+\frac{\omega(t)-\mu_{\omega}}{\mu_{\omega}}-\gamma\right\},&t>0;\\ \omega(t),&t=0;\end{cases}$ (2)
@@
$t_{E_{x}}=\operatorname*{arg\,max}_{t\in[t_{\text{start}},t_{\text{end}}]}\Gamma(E_{x},t)$ (3)
@@
$\Gamma(E_{y},t_{E_{y}})>\Gamma(E_{x},t_{E_{x}}),\;\forall\,x\in\mathcal{N}\,\backslash\,\mathcal{N}_{E}.$ (5)
@@
$EI_{x}=\frac{c_{E_{x}}}{t_{E_{x}}-t_{\text{start}}},$ (6)
@@
3.2 Effective Connectivity Model
To estimate the SEEG connectivity dynamics, we consider the PTE algorithm, which, given a couple of channels $x,y\in\mathcal{N}$ at time $t$, models their coupling according to the phase distributions of $x(t)$ and $y(t)$. From this perspective, PTE may lead to different results depending on how we model the phase distribution. Given a specific window $x(t)$, we first compute its analytic representation $\mathcal{X}(t)=x(t)+HT\left(x(t)\right)$, where $HT(\cdot)$ is the Hilbert transform. Notably, $\mathcal{X}(t)$ is associated with $n=T_{\text{window}}/f_{s}$ phase values, representing the evolution of the instantaneous phase $\theta_{x}(t)$ of $x(t)$ in time. According with the Sturges rule [41], we can infer the bin width $\vartheta$ of $\theta_{x}(t)$ as $\vartheta=2\pi/(\log_{2}(n)+1)$.
The proposed methodology quantizes the phase values to reduce the influence of volume conduction on the connectivity model. A similar same principle is adopted by the PLI algorithm, which, however, does not consider the effectiveness of the network connection [19]. In addition, PLI presents critical pitfalls, including underestimating those connections associated with small propagation delays or with signals whose phase difference distribution fluctuates around zero. Hence, PTE tries to act as a compromise between PLI and PLV, which, notably, suffers the effect of volume conduction and leads to many false connectivity phenomena.
Lets denote by $\theta_{x}(t)$ and $\theta_{y}(t)$ the phase distribution of $x(t)$ and $y(t)$, respectively. According to the PTE algorithm, the connectivity that the channel $x$ exerts on the channel $y$ at time $t$, considering a lag $\tau$, is given by
$\begin{gathered}T_{x\rightarrow y}(t,\tau)=H\left(\theta_{y}(t),\theta_{y}(t+\tau)\right)+\\ H\left(\theta_{x}(t),\theta_{y}(t)\right)-H\left(\theta_{x}(t),\theta_{y}(t),\theta_{y}(t+\tau)\right)-H\left(\theta_{y}(t)\right),\end{gathered}$ (7)
where $H(\cdot)$ denotes the entropy function. The PTE formulation, as reported in (7), is a directed connectivity measure, which means that, in general, $T_{x\rightarrow y}(t,\tau)\neq T_{y\rightarrow x}(t,\tau)$. In order to remove the dependency from $\tau$, with a slight abuse of notation, we redefine the phase transfer entropy between $x(t)$ and $y(t)$ as the maximum value of $T_{x\rightarrow y}(t,\tau)$ among multiple lags in the set $[0,\tau_{max}]$:
$T_{x\rightarrow y}(t)=\max_{\tau\in[0,\tau_{max}]}T_{x\rightarrow y}(t,\tau).$ (8)

### v4

@@
Abstract.
In this work, we propose a new computational framework to assist neurophysiologists in Stereoelectroencshalography (SEEG) analysis, with the final aim of improving the definition of the Epileptogenic Zone (EZ) in patients with drug-resistant epilepsy. We design a new algorithm, named Desynchronization Index (DI), that classifies as most epileptogenic those SEEG channels that show independent behavior during the seconds preceding the seizure propagation. We test the proposed DI algorithm against the Epileptogenic Index (EI) algorithm on a clinical dataset of 11 patients, considering the neurophysiological evaluation of the EZ as the clinical ground truth. Our results denote that DI overcomes EI in terms of area under the ROC curve (AUC=0.81 vs AUC=0.74) while combining the two algorithms as a unique tool leads to the best performance (AUC=0.87). The DI algorithm underscores connectivity dynamics that can hardly be identified with a pure visual analysis, increasing the accuracy of the EZ definition compared to traditional methods. This technique can lead to the definition of a new effective biomarker of the EZ, reducing the burden required by the SEEG review in the case of extensive implants and improving our understanding of the dynamics leading to the generation of seizures.
@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXjournal: HEALTHjournalvolume: 1journalnumber: 1article: 1publicationmonth: 1ccs: Applied computing Health informaticsccs: Theory of computation Dynamic graph algorithms
1. Introduction
@@
The localization of the Epileptogenic Zone (EZ) is the fundamental prerequisite for performing epilepsy surgery in people with drug-resistant epilepsy (Lüders et al., 2006). Ensuring a higher accuracy in the definition of the EZ makes it possible to reduce the extent of the surgery, mitigating the risk of complications or chronic deficits. In the most complex scenario, the EZ can be identified only through invasive procedures, among which Stereoelectroencephalography (SEEG) represents the most accurate methodology (Isnard et al., 2018). An SEEG recording involves the surgical implantation of electrodes into the patient’s brain, enabling the monitoring of electrical activity from deep cortical regions, which is essential for the identification of the EZ. As a result, SEEG-guided surgical resections lead more than $60\%$ of patients to be seizure-free after the intervention (Cossu et al., 2005). However, SEEG has the drawback of providing very focused information: if none of the intracranial electrodes intersects the EZ, the procedure may result in an unclear EZ definition and an unsuccessful epilepsy surgery (Cardinale et al., 2013).
@@
Nowadays, SEEG interpretation is mainly based on visual analysis, focusing on the early phase of the seizure discharge. The final goal is to define an EZ that can be focal or distributed among multiple cortical structures, a scenario falling under the concept of epileptogenic networks (Bartolomei et al., 2017a). The above process is particularly complex since SEEG signals include hundreds of components, whose analysis requires the involvement of highly specialized neurophysiologists. For this reason, in the last years, SEEG interpretation has started being supported by computational tools that aim at characterizing the EZ and, more in general, the dynamics leading to seizure generation (Gnatkovsky et al., 2014). The most common methods analyze the SEEG power spectrum, searching for the cortical sites generating fast oscillations (Roehri et al., 2016), i.e., electrical activities with frequencies in the gamma range ($[30,100]$ Hz), as done in the cases of Epileptogenic Index (EI) (Bartolomei et al., 2008) and Epileptogenic Maps (EMs) (David et al., 2011).
@@
Another approach is to model cortical connectivity with the aim of identifying which brain structures show abnormal connectivity behavior during seizure onset. In this context, the first studies analyzed the responses generated through intracranial stimulations, known as Cortico-Cortical Evoked Potentials (CCEPs) (Bhattacharyya et al., 2019), which, however, do not have a physiological origin and may not correctly represent cortical connectivity patterns. Other solutions estimate the exchange of information by analyzing the phase distribution of the signals associated with each cortical site, as performed by the Phase Locking Value (PLV) and the Phase Lag Index (PLI). These methods avoid biases associated with amplitude differences within the signal recordings but only allow us to estimate the synchrony between different signals, which results in a functional connectivity model. To estimate the direction of neural connections, it is necessary to use more advanced methodologies such as the Phase Transfer Entropy (PTE) which, however, was applied to SEEG data only in a limited number of cases.
@@
Although multiple quantitative approaches have been developed for the study of epileptogenic networks, the literature often presents discordant findings. Most works agree that the ictal phase (i.e., the period that involves the epileptic discharge) is preceded by a reduction in cortical connectivity, followed by the opposite phenomenon during the seizure propagation (Proix et al., 2017). Recently, it has been hypothesized that the EZ presents an abnormal inward information flow during the inter-ictal phase (Gunnarsdottir et al., 2022), and the epileptogenic level of each cortical site can be assessed by the connection exerted during seizures (Nithin et al., 2021).
@@
In this work, we present a new framework for supporting neurophysiologists in SEEG interpretation. Specifically, we design a connectivity-based algorithm, named Desynchronization Index (DI), that exploits the PTE to estimate the relations between brain structures and detect connectivity anomalies in the resulting time-varying network. According to our approach, the EZ is identified by the cortical sites that present an independent behavior during the ictal phase. Overall, the main contributions of this manuscript consist of the following points:
@@ 2. Related Work
In clinical practice, the interpretation of SEEG recordings is mainly based on visual analysis and constitutes a highly time-consuming process because of the huge number of data that must be reviewed (Bartolomei et al., 2017b). The most common marker of the seizure onset is given by low-voltage fast oscillations, which are electrical discharges spanning beta frequencies ($[13,30]$ Hz), usually observed in mesial temporal seizures, and gamma frequencies ($[30,100]$ Hz), observed in neocortical seizures. To quantify the magnitude of fast oscillations and the network dynamics leading to seizures, it is possible to analyze the spectrum of SEEG signals. This approach is followed by the EI algorithm (Bartolomei et al., 2008), which is the most routinely used method to assist clinicians in SEEG interpretation (Makhalova et al., 2023).
@@
The EI algorithm, occasionally combined with other biomarkers (Balatskaya et al., 2020), has been recognized to provide the best accuracy in terms of EZ localization and surgical prediction. An extension of such an approach is the epileptogenicity rank, a technique that, before computing the EI values, assigns weights to SEEG sites depending on to their distance from the hypothesized center of the EZ (Parasuram et al., 2021). This solution has the obvious limitation of biasing the epileptogenic levels assigned to each channel by a prior assumption. In particular, the epileptogenicity rank framework reduces the number of false positives when the clinical hypothesis is correct, but it dramatically decreases sensitivity in the other case.
@@
While EI focus on the energy distribution in the frequency domain, other approaches analyze the variation in neuronal connectivity during the ictal phase. Several tools have been proposed to this goal, including the PLV (Lachaux et al., 1999) and the PLI (Stam et al., 2007), with different trade-offs in terms of sensitivity and precision (Gupta et al., 2020). In particular, the PLI was proposed as an alternative to the PLV in order to mitigate the noisy information associated with the volume conduction (Peraza et al., 2012), reducing the number of false positives. A more advanced method is the PTE (Lobier et al., 2014), which exploits the Granger Casuality (GC) (Shojaie and Fox, 2022) to estimate both the delay and the direction of neuronal connections, and constitutes a very promising technique for SEEG analysis (Wang et al., 2017).
@@
In addition to designing EI, Bartolomei et al. analyze the SEEG data using the nonlinear regression coefficient (h2) (Wendling et al., 2000), and show that the epileptogenic network is associated with reduced connectivity in the period that anticipates the ictal phase  (Bartolomei et al., 2001). In particular, the h2 coefficient is agnostic to the signal frequencies, while most connectivity models discern cortical relationships depending on the frequency bands. Following this principle, the authors of (Daoud et al., 2022) estimate connectivity by the cross-correlation between SEEG signals, showing that seizure frequency is proportional to the outward connectivity of epileptogenic channels. Similar methodologies are given in (Bernabei et al., 2021), where pairwise correlation is combined with Euclidean distance to estimate the EZ extension, and in (Li et al., 2021), where graph theory is exploited to identify the network hubs that contribute the most to seizure propagation.
@@
Recently, the use of Spearman rank correlation (Zar, 2005) to assess the coupling between SEEG signals has shown good results for both the detection of the EZ and the prediction of surgical outcomes (Huang et al., 2023). Other connectivity-based approaches associate the phase of slow oscillations ($<30$ Hz) with the amplitude of fast oscillations ($>30$ Hz), as occurred for the Phase Slope Index (PSI) and the Phase Amplitude Coupling (PAC) (Bastos and Schoffelen, 2016). In a recent work (Jiang et al., 2022), the use of PSI allows the authors to show that the EZ presents higher outward connectivity during seizures than in the resting state. Similar results are observed in (Liu et al., 2021), where SEEG connectivity is estimated via the PAC between the high (gamma) and low (delta and theta) frequency bands. The authors of (An et al., 2020) seem to contradict these findings and, using the Partial Direct Coherence (PDC) (Baccalá and Sameshima, 2001) as the connectivity model, assess that the EZ presents higher inward connectivity during seizure generation.
@@
In recent years, the popularity of Machine Learning (ML) has encouraged many researchers to apply data-driven techniques to epileptogenic networks (Hussein et al., 2020). For example, the authors of (Guo et al., 2021) exploit hypergraph learning to automatically detect SEEG fast oscillations, while (Hu et al., 2023) investigates different patient-specific ML models for analyzing scalp electroencephalography signals. In addition, (Grinenko et al., 2017) exploit a Support Vector Machine (SVM) model to integrate multiple electrophysiological characteristics, including both fast oscillations and low-frequency suppression, to improve accuracy in EZ detection. This supports the hypothesis that ictal channels are not identified by specific frequency components, and more comprehensive methodologies are required to describe the generation of seizures. On the other hand, (Grinenko et al., 2017) and similar works consider as ground truth the area resected by epilepsy surgery, which is usually much larger than the EZ to maximize the probability of seizure reduction. In general, the high complexity behind SEEG analysis makes it extremely difficult to build supervised datasets, which prevents ML solutions from obtaining reliable performance in this field (Wu et al., 2024).
@@
Although the large variety of works implementing connectivity-based metrics for the interpretation of SEEG, the results of the literature do not lead to an agreement on which biomarkers characterize EZ. Initial studies show that the ictal transition is characterized by a reduction in connectivity, but the EZ is often identified by cortical sites with a greater coupling during seizure propagation. In this work, we try to answer these questions by quantifying the level of epileptogenicity of SEEG sites in terms of desynchronization. With the latter, we denote the tendency of a cortical site to show independent electrical behavior that cannot be inhibited by neighboring brain structures.
3. Analytical Model
@@
In this section, we first describe how SEEG signals are modeled within our framework and introduce the EI algorithm, which represents the state-of-the-art for detecting the EZ. Then, we present the connectivity model used to infer the relation between SEEG sites starting from their phase distribution. Finally, we introduce the DI algorithm, which evaluates the epileptogenicity level of each site according to its tendency to isolate itself from the network.
3.1. Epileptogenic Index
@@
We model a single SEEG recording as a multidimensional signal with $N$ different components, named channels, one for each cortical site analyzed. In the following, we denote by $\mathcal{N}$ the set of channels and by ||$\cdot$|| ⋅ | the cardinality operator, so that the number of cortical sites is equal to $N=|\mathcal{N}|$. According to our approach, each channel $x\in\mathcal{N}$ is segmented in multiple overlapping windows $x(t)$, with $t=0,\Delta t,2\Delta t,\text{...}$, where $\Delta t$ represents the time-shift between consecutive windows. Given the window duration $T_{\text{window}}$ and the sampling frequency $f_{s}$, the sample number per window is $n=T_{\text{window}}/f_{s}$.
@@
The EI algorithm requires to evaluate the signal energy in the frequency domain (Bartolomei et al., 2008). Given a specific window $x(t)$, the algorithm first computes the Fourier Transform (FT) of $x(t)$, obtaining a complex value $X(t,f)$ for each frequency in $[0,f_{s}/2]$. Hence, the energy ratio between the high- and low-frequency bands of the target signal $x$ is computed as
@@
(2) $\Gamma(\omega,t)=\begin{cases}\max\left\{0,\Gamma(\omega,t-\Delta t)+\frac{\omega(t)-\mu_{\omega}}{\mu_{\omega}}-\gamma\right\},&t>0;\\ \omega(t),&t=0;\end{cases}$
@@
(3) $t_{E_{x}}=\operatorname*{arg\,max}_{t\in[t_{\text{start}},t_{\text{end}}]}\Gamma(E_{x},t)$
@@
(5) $\Gamma(E_{y},t_{E_{y}})>\Gamma(E_{x},t_{E_{x}}),\;\forall\,x\in\mathcal{N}\,\backslash\,\mathcal{N}_{E}.$
@@
(6) $EI_{x}=\frac{c_{E_{x}}}{t_{E_{x}}-t_{\text{start}}},$
@@
3.2. Effective Connectivity Model
To estimate the information flows within the SEEG network, we consider the PTE algorithm, which, given a couple of channels $x,y\in\mathcal{N}$ at time $t$, models their coupling according to their phase distributions. Given a specific window $x(t)$, we first compute its analytic representation $\mathcal{X}(t)=x(t)+HT\left(x(t)\right)$, where $HT(\cdot)$ is the Hilbert transform. We observe that $\mathcal{X}(t)$ is associated with $n=T_{\text{window}}/f_{s}$ phase values, representing the evolution of the instantaneous phase $\theta_{x}(t)$ of $x(t)$ in time.
As next step, we need to model the phase distribution: in this work, we follow the Sturges rule (Freedman and Diaconis, 1981) and compute the bin width of $\theta_{x}(t)$ as $\vartheta=2\pi/(\log_{2}(n)+1)$. This quantization process allows us to reduce the influence of volume conduction on the connectivity model. A similar principle is adopted by the PLI, which, however, does not consider the direction of the information flows between neural signals (Stam et al., 2007). In addition, PLI underestimates the neural connections associated with small propagation delays or whose phase difference distribution fluctuates around zero. The opposite approach is taken by the PLV, which model the phase in a continuous domain and, consequently, suffers more the problems associated with volume conduction. By varying the number of phase bins in our model, we can trade off between specificity and sensitivity, similar to what occurs when using the PLI instead of the PLV.
Let us denote by $\theta_{x}(t)$ and $\theta_{y}(t)$ the phase distribution of $x(t)$ and $y(t)$, respectively. According to the PTE algorithm, the connectivity that the channel $x$ exerts on the channel $y$ at time $t$, considering a lag $\tau$, is given by
(7) $\begin{gathered}T_{x\rightarrow y}(t,\tau)=H\left(\theta_{y}(t),\theta_{y}(t+\tau)\right)+\\ H\left(\theta_{x}(t),\theta_{y}(t)\right)-H\left(\theta_{x}(t),\theta_{y}(t),\theta_{y}(t+\tau)\right)-H\left(\theta_{y}(t)\right),\end{gathered}$
where $H(\cdot)$ denotes the entropy function. We observe that the PTE, as reported in (7), is a directed connectivity measure, which means that, in general, $T_{x\rightarrow y}(t,\tau)\neq T_{y\rightarrow x}(t,\tau)$. In order to remove the dependency from $\tau$, with a slight abuse of notation, we redefine the phase transfer entropy between $x(t)$ and $y(t)$ as the maximum value of $T_{x\rightarrow y}(t,\tau)$ among multiple lags in the set $[0,\tau_{max}]$:
(8) $T_{x\rightarrow y}(t)=\max_{\tau\in[0,\tau_{max}]}T_{x\rightarrow y}(t,\tau).$